# Predicting New Research Directions in Materials Science using Large Language Models and Concept Graphs

### v1

@@ Results
An overview of the report generation is shown in Figure 7. First, a set of individualized concepts $C_{{\text{own}}}$ is generated as the intersection of (i) all concepts extracted from the abstracts of the recent publications of the respective researcher and (ii) all known concepts $C_{{\text{known}}}$ in the concept graph. Based on these two sets of concepts, we generated researcher-specific suggestions, i.e. combinations of concepts (SI Section Report generation): The first two report sections, $S_{\text{own}\times\text{own}}$ and $S_{\text{own}\times\text{other}}$, contain the top 25 combinations of the own concepts with themselves and with all other concepts, respectively. We applied two heuristics (avoid generic concepts and avoid too similar as well as too unrelated cobinations based on their semanitic embeddings) to filter the suggestions in the second category, resulting in a section $S^{\text{filtered}}_{\text{own}\times\text{other}}$. To take the researcher’s full profile into account, the next section $S_{\text{(many own)}\times\text{other}}$ contains the top 20 concepts with highly scored connections to many own concepts. For the final section of the report (‘LLM curation’), LLMs were queried to select interesting combinations from the previous sets of combinations and to write a short paragraph with more information on how the concepts can be combined and why these specific combinations are promising new research directions.
@@
Based on the individual reports described above, a 30-minute interview was conducted with each of the researchers, in which the suggested combinations of topics were classified as already known (A), nonsensical or not understandable (B), and novel, interesting, or inspiring (C). For further analysis, the first class was further divided into already published combinations (A1) which were likely missed during dataset generation (e.g., very recent publications or publications outside of the analyzed literature corpus) and obvious, trivial, or very general combinations (A2), which are not necessarily mentioned together in an abstract.
The occurrence of the combinations classified as A1 (known), A2 (trivial), B (nonsense), and C (interesting) using this scheme is listed in Table 3. Out of 292 categorized suggestions, we find the following: Class A1: 71 suggestions, Class A2: 36 suggestions, Class B: 99 suggestions, and Class C: 86 suggestions. An excerpt of combinations per category can be found in Table LABEL:si-tab:ranked-combinations in the SI. As mentioned above, the number of interview partners is not sufficient for a reliable statistical analysis, e.g. the total number of classifications per researcher ranged between 18 and 48. An overview of the classified combinations per researcher can be found in the SI (Figure S8). During the interviews, the participants were sometimes unsure whether to label a suggestion as B or C. To present a clearer overview, the uncertain classifications were counted as C, as the suggestion was at least partly interesting or inspiring. In addition, the concepts in $C_{\text{(many own)}\times\text{other}}$ were also generously counted as an overall interesting suggestion (category C), i.e. already if one of the many concepts in $C_{\text{own}}$ was inspiring in conjunction with the proposed concept.
@@
C 29 33.0%
To evaluate the usefulness of the ‘LLM Curation’ approach, we analyze how many of the combinations suggested by an LLM are later labeled as interesting. In Table 4, we display the confusion matrix of the variables "Suggested by an LLM" and "Is Interesting", i.e. whether the suggestion was rated as C by a human expert. Interestingly, we observe a rounded precision of 51% regarding the LLM selecting interesting concepts, i.e. 27 of the 52 concepts suggested by the LLM as interesting are also labeled as interesting by the scientists. This is a substantial improvement in precision compared to 70 of 266 (26%) which were analyzed in total. The recall is not of primary interest in this context as the LLM was only allowed to select a limited number of combinations as described above. An analysis of the previous node distances of the suggested combinations across all reports showed that 6 out of 9 concept pairs at $d_{\text{prev}}=3$, were rated as category ‘C’. This high ratio of interesting combinations underpins our previous assumption that the inclusion of semantic information in the prediction models increases its capability to foster out-of-the-box thinking.
Table 4: LLM suggestion vs. "Is Interesting" (labeled as C) confusion matrix. The suggested combinations from section $S_{\text{(many own)}\times\text{other}}$ are not present as the ‘LLM Curation’ did not comprise those combinations.
LLM suggestion
Y N $\sum$
"Is Interesting" Y 27 43 70
N 26 170 196
$\sum$ 53 213 266
@@
To illustrate the value of the suggested combinations of concepts in more detail, we discuss below five selected examples of category C suggestions by putting them into context and describing in more detail why these combinations are interesting. More detailed discussions of all five topics can be found in the Supporting Information.
@@
This suggestion associates "conventional ceramics" with "graphene oxide", two domains seldom combined. Conventional oxide ceramics provide chemical, thermal, and structural stability. Graphene oxide offers a high-surface-area, electronically conductive carbon framework. Their union could yield composites marrying ceramic robustness with rapid charge and heat transport, relevant to batteries, catalysis, and thermal barriers. Existing studies mix pre-synthesised oxides with graphene derivatives, giving limited interfacial contact. Our unpublished autoclave route aligns with the AI hint: Fe(CO)5 decomposes on graphene oxide, forming a $\approx$200 nm iron-oxide shell on multilayer graphene (see Figure 8. The in-situ process creates intimate oxide–graphene interfaces and a continuous conductive network. Electrochemical tests show high reversible capacity and enhanced redox kinetics in Li-ion conversion cells. These findings demonstrate the model’s capacity to provide inspiration for overlooked but feasible synthesis strategies. Systematic exploration of AI-highlighted ceramic/graphene hybrids may accelerate multifunctional material discovery.
[…]
@@
Thin-film organic and perovskite solar cells comprise multilayers with mismatched thermal-expansion coefficients. Temperature excursions during coating, annealing, and operation impose tensile strain at these organic–inorganic interfaces. Such strain drives delamination and point-defect formation, accelerating performance loss. While strain engineering is routine in inorganic semiconductors, it is rarely applied in soft-matter photovoltaics. Molecular architecture provides a complementary lever: greater torsional flexibility lowers film modulus and dissipates stress. Thus, the AI-proposed link “tensile strain + molecular architecture” highlights an under-exploited stability pathway. Brabec & Friederich (Science 2024) showed that TPA-containing hole-transport layers benefit efficiency by accommodating strain. An unpublished follow-up finds plasticizing additives further boost mechanical and operational durability. These studies corroborate strain-aware molecular design as a broadly applicable interface strategy. Systematic exploration could extend device lifetimes without compromising performance.
@@
Ferroelectric in-plane polarization, recently demonstrated in MAPbI3 perovskites, spatially separates photocarriers and channels them toward electrodes. The ML suggestion indicates that similar lateral dipole fields could be engineered in organic absorbers. Asymmetric polar moieties, oriented during self-assembly or within covalent organic frameworks, may supply the required non-centrosymmetry. The resulting internal field should enhance carrier separation and transport, while a higher dielectric constant lowers exciton binding and monomolecular recombination. Ferroelectricity is so far verified only for halide perovskites [rossi_ferroelectric_2018, rohm_ferroelectric_2017, schulz_ferroelectricity_2022], and is absent in silicon or conventional organic cells [rohm_ferroelectric_2019, breternitz_role_2020]. Piezoelectric polymers such as PVDF already exploit oriented dipoles in sensing devices, suggesting viable processing routes. Prior attempts to raise organic permittivity show limited success, and deliberate in-plane polarization remains unexplored [roehm2020]. Hence, the predicted concept defines a tractable, novel direction for photovoltaic materials research.
@@ Discussion
In the first part of this study, we showed that the power of large language models, especially LLaMa-2-13B, can be harnessed to extract scientific concepts – vaguely defined as key phrases – from scientific text. We established a methodology for fine-tuning open-source LLMs based on a small manually labeled abstracts, which guides the LLM to extract only relevant concepts. The initial training data can be iteratively extended by humanly corrected LLM annotations to further improve the extraction process, but no human verification is required to check the final 221,000 labeled data points. Follow-up studies may investigate whether prioritizing quality over quantity [zhou_lima_2023] in the annotated training examples, i.e. by using fewer but carefully selected data points, could yield more accurate and representative extracted concepts, and also whether the inclusion of synthetic data can help to accelerate the annotation process and further enhance model performance.
To achieve the second goal of this work, we created a concept graph, derived from the previously extracted materials science concepts and the dates of the corresponding publications. This graph was then successfully used to predict emerging links between previously unconnected concepts, underscoring that a simple graph representation suffices for this task. Finally, we demonstrated that the integration of semantic knowledge in the form of concept embeddings boosts the predictive performance of our model. The usefulness of our model in a real-world scenario was investigated through qualitative interviews with domain experts.

### v2

@@ Results
An overview of the report generation is shown in 3. First, a set of individualized concepts $C_{{\text{own}}}$ is generated as the intersection of (i) all concepts extracted from the abstracts of the recent publications of the respective researcher and (ii) all known concepts $C_{{\text{known}}}$ in the concept graph. Based on these two sets of concepts, we generated researcher-specific suggestions, i.e. combinations of concepts: The first two report sections, $S_{\text{own}\times\text{own}}$ and $S_{\text{own}\times\text{other}}$, contain the top 25 combinations of the own concepts with themselves and with all other concepts, respectively. We applied two heuristics (avoid generic concepts and avoid too similar as well as too unrelated combinations based on their semantic embeddings) to filter the suggestions in the second category, resulting in a section $S^{\text{filtered}}_{\text{own}\times\text{other}}$. To take the researcher’s full profile into account, the next section $S_{\text{(many own)}\times\text{other}}$ contains the top 20 concepts with highly scored connections to many own concepts. For the final section of the report (‘LLM curation’), LLMs were queried to select interesting combinations from the previous sets of combinations and to write a short paragraph with more information on how the concepts can be combined and why these specific combinations are promising new research directions. A technical definition of each section is given in section S4.
@@
Based on the individual reports described above, a 30-minute interview was conducted with each of the researchers, in which the suggested combinations of concepts were classified as already known (A), nonsensical or not understandable (B), and novel, interesting, or inspiring (C). In section 4 ($S_{\text{(many own)}\times\text{other}}$), suggestions were generously already counted as overall interesting (category C) if one of the many concepts in $C_{\text{own}}$ was inspiring in conjunction with the proposed other concept. To account for cases in which the participants were unsure whether to label a suggestion as B or C, an additional category D was introduced in the analysis. For further analysis, the first class was divided into already published combinations (A1) which were likely missed during dataset generation (e.g., very recent publications or publications outside of the analyzed literature corpus) as well as obvious, trivial, or very general combinations (A2), which are not necessarily mentioned together in an abstract.
The occurrence of the combinations classified as A1 (known), A2 (trivial), B (nonsense), C (interesting), D (uncertain) using this scheme is listed in 3. Out of 292 categorized suggestions, we find the following: Class A1: 71 suggestions, Class A2: 36 suggestions, Class B: 99 suggestions, Class C: 77 and Class D: 9, suggestions. Thus, 26% of all suggested concept combinations were considered interesting by the interviewees. An excerpt of combinations per category can be found in Supplementary Table 7 and Supplementary Table 8. As mentioned above, the number of interview partners is not sufficient for a reliable statistical analysis, e.g. the total number of classifications per researcher ranged between 18 and 48, with a per-participant variance ranging from 5.04 to 51.36. An overview of the classified combinations per researcher and the per-participant variance can be found in (Supplementary Fig. 20 and Supplementary Table 6.
@@
C 26 29.5%
D 3 3.4%
To evaluate the usefulness of the ‘LLM Curation’ approach, we analyze how many of the combinations suggested by an LLM are later labeled as interesting. In Supplementary Table 5, we display the confusion matrix of the variables "Suggested by an LLM" and "Is Interesting", i.e. whether the suggestion was rated as C by a human expert. Interestingly, we observe a rounded precision of 47% regarding the LLM selecting interesting concepts, i.e. 24 of the 53 concepts suggested by the LLM as interesting were also labeled as interesting by the scientists. This is a substantial improvement in precision compared to 61 of 266 (23 %) concept combinations analyzed in total. In this context the recall is not of primary interest, as the LLM was only allowed to select a limited number of combinations as described above. An analysis of the previous node distances of the suggested combinations across all reports showed that 5 out of 9 concept pairs at $d_{\text{prev}}=3$, were rated as category ‘C’. This high ratio of interesting combinations underpins our previous assumption that the inclusion of semantic information in the prediction models increases its capability to foster out-of-the-box thinking.
@@
To illustrate the value of the suggested combinations of concepts in more detail, we discuss below five selected examples of category C suggestions by putting them into context and describing why these combinations are interesting. A more detailed discussion of all five concept combinations can be found in the Supporting Information.
Figure 4: Exfoliated graphene oxide (resulting in multilayer graphene) covered with $200\text{\,}\mathrm{n}\mathrm{m}$ thick iron oxide layers. a) shows the flake morphology of exfoliated layers, b) shows the iron oxide layer covering the whole surface.
@@
This suggestion associates "conventional ceramics" with "graphene oxide", two domains seldom combined. Conventional oxide ceramics provide chemical, thermal, and structural stability. Graphene oxide offers a high-surface-area, electronically conductive carbon framework. Their union could yield composites marrying ceramic robustness with rapid charge and heat transport, relevant to batteries, catalysis, and thermal barriers. Existing studies mix pre-synthesised oxides with graphene derivatives, giving limited interfacial contact. We show preliminary data of a $\approx 200$ nm iron-oxide shell on multilayer graphene in 4 (unpublished). The in-situ process creates intimate oxide–graphene interfaces and a continuous conductive network. Electrochemical tests show high reversible capacity and enhanced redox kinetics in Li-ion conversion cells. These findings demonstrate the model’s capacity to provide inspiration for overlooked but feasible synthesis strategies. Systematic exploration of AI-highlighted ceramic/graphene hybrids may accelerate multifunctional material discovery.
@@
Thin-film organic and perovskite solar cells comprise multilayers with mismatched thermal-expansion coefficients. Temperature excursions during coating, annealing, and operation impose tensile strain at these organic–inorganic interfaces. Such strain drives delamination and point-defect formation, accelerating performance loss. While strain engineering is routine in inorganic semiconductors, it is rarely applied in soft-matter photovoltaics. Molecular architecture provides a complementary lever: greater torsional flexibility lowers film modulus and dissipates stress. Thus, the AI-proposed link “tensile strain + molecular architecture” highlights an under-exploited stability pathway. In 2024, it was shown by Brabec & Friederich that TPA-containing hole-transport layers benefit efficiency by accommodating strain[31]. These studies corroborate strain-aware molecular design as a broadly applicable interface strategy. Systematic exploration could extend device lifetimes without compromising performance.
@@
Ferroelectric in-plane polarization, recently demonstrated in MAPbI3 perovskites, spatially separates photocarriers and channels them toward electrodes. The ML suggestion indicates that similar lateral dipole fields could be engineered in organic absorbers. Asymmetric polar moieties, oriented during self-assembly or within covalent organic frameworks, may supply the required non-centrosymmetry. The resulting internal field should enhance carrier separation and transport, while a higher dielectric constant lowers exciton binding and monomolecular recombination. Ferroelectricity is so far verified only for halide perovskites [32, 33, 34], and is absent in silicon or conventional organic cells [35, 36]. Piezoelectric polymers such as PVDF already exploit oriented dipoles in sensing devices, suggesting viable processing routes. Prior attempts to raise organic permittivity show limited success, and deliberate in-plane polarization remains unexplored [37]. Hence, the predicted concept defines a tractable, novel direction for photovoltaic materials research.
@@ Discussion
In the first part of this study, we showed that the power of large language models, especially LLaMa-2-13B, can be harnessed to extract scientific concepts – vaguely defined as key phrases – from scientific text. We established a methodology for fine-tuning open-source LLMs based on a small manually labeled abstracts, which guides the LLM to extract only relevant concepts. The initial training data can be iteratively extended by humanly corrected LLM annotations to further improve the extraction process, but no human verification is required to check the final 221,000 labeled data points. Follow-up studies may investigate whether prioritizing quality over quantity [38] in the annotated training examples, i.e. by using fewer but carefully selected data points, could yield more accurate and representative extracted concepts, and also whether the inclusion of synthetic data can help to accelerate the annotation process and further enhance model performance.
To achieve the second goal of this work, we created a concept graph, derived from the previously extracted materials science concepts and the dates of the corresponding publications. This graph was then successfully used to predict emerging links between previously unconnected concepts, underscoring that a simple graph representation suffices for this task. Finally, we demonstrated that the integration of semantic knowledge in the form of concept embeddings boosts the predictive performance of our model. Combining the GNN approach with semantic features is possible and will be explored in future work. The usefulness of our model in a real-world scenario was investigated through qualitative interviews with domain experts, who rated 77 or 26% out of 292 generated recommendations as interesting. While this rate may sound modest, each 30-minute session still yielded several promising ideas, making the outcome practical for guiding research.
@@
Methods
The key steps of our approach are depicted in 5. After gathering the abstracts of a large number of research publications in the domain of materials sciences, we extracted the main concepts, i.e. short key phrases consisting of few words from these abstracts, and used them as nodes in a concept graph that mirrors the (time-dependent) connectivity of the materials science concepts in literature. The final step of our workflow consists of performing link prediction on this graph based on both network properties, e.g. connectivity information, and semantic knowledge about the concepts captured in aggregated word embeddings.
Figure 5: Overview of the link prediction workflow: (a) gathering of materials science abstracts with OpenAlex [39], (b) extraction of concepts using LLMs, (c) creation of the semantics-aware-concept graph, and (d) prediction of new research directions.
Dataset
We prepared a dataset of published papers related to materials science. Data was obtained from OpenAlex by querying all publications listed at materials science-related journals, conferences, and other venues [39]. The retrieved papers were filtered based on language, length, and whether they have an abstract. For each publication, the title and abstract were cleaned and concatenated. Chemical formulae were extracted, stored separately, and later merged with the extracted concepts. The resulting dataset comprises approximately 221,000 articles published between 1955 and 2022, with relevant attributes being "title", "abstract", and "publication date". A more detailed description of the dataset generation is given in S7.
Concept extraction
Previous work used RAKE for concept extraction in conjunction with manual filtering to remove errors, i.e. phrases that do not represent semantic information, introduced by the imperfect statistical analysis performed in RAKE [12, 14, 13]. Instead, we opted for extracting concepts using fine-tuned LLMs (see Fig. 6). To create a dataset for fine-tuning, 100 randomly chosen abstracts were first manually annotated by extracting and partially adjusting or even paraphrasing relevant and meaningful concepts as a preliminary step. We note that manual annotation is particularly sensitive to the labeler, as there is no unique way of extracting and defining concepts. Subsequently, we fine-tuned the LLaMa-2-13B base model [40, 41] on our manually annotated abstracts for 4 epochs, using a learning rate of $5\cdot 10^{-4}$ and a batch size of 1 (see S9). Llama-2 models were SOTA when the work was performed in 2023, but will be replaced with newer models in future iterations of this work). The size of the model is a trade-off between accuracy and cost, as it is the largest model that can process 20 abstracts at once on an A100 GPU with 80GB of VRAM. To accelerate training and especially inference, we incorporated 8-bit quantization [42] and low-rank adaptation (LoRa) techniques [43, 44] using Hugging Face’s PEFT module [45].
Similar to Dunn et al.’s assisted annotation process [4], the fine-tuned model’s outputs were compared in the third step to the concepts extracted by GPT-3.5 [7] to efficiently identify and correct common mistakes of the fine-tuned model. In this way, we labeled 100 additional abstracts, and the base model was again fine-tuned on the larger dataset of 200 abstracts. The process of iteratively adding more automatically extracted and manually corrected concepts to fine-tune the model with more datapoints could, in principle, be repeated more often, but 200 labeled abstracts were enough for our use case. Finally, the resulting model was employed to extract concepts from all approximately 221,000 abstracts in our dataset, requiring approximately 160 GPU-hours. Future updates of our concept graph are substantially less demanding, as only incremental (delta) extraction is required. Future developments in LLM research might enable a complete re-evaluation with higher quality and reliability. After extraction, we conducted minor post-processing of the extracted concepts by removing the remaining plural forms. We note that some bias was introduced by the selection of 200 abstracts with 3102 distinct concepts, as they were all from materials science, and observed that, e.g., core-biology concepts were not extracted by our method.
Concept graph
To enrich the nodes of our concept graph with semantic information, we calculate concept embeddings and use them as node features. The procedure of calculating concept embeddings using MatSciBERT [17] is summarized in Fig. 7: First, both the entire abstract and the previously extracted concepts are tokenized, and the embeddings are then calculated for the tokenized abstract. The next step consists of locating all instances of a concept in the abstract and averaging the embeddings of the tokens corresponding to the concept. For example, the concept ‘mechanical stress’ is tokenized as [4487, 1893], and the concept embedding is calculated as the average of the corresponding representations in the embedded abstract, i.e. those corresponding to the positions of the sequence [4487, 1893] in the tokenized abstract. To derive a singular representation for each concept per abstract, we average the embeddings of all its occurrences. In cases where a concept does not appear verbatim in an abstract — e.g. due to the normalization processes during the initial concept extraction — we take the mean embedding of all tokens (while excluding the start and end token in the abstract as its representation). As the final step, the average embedding for identical concepts across different abstracts is calculated to obtain a single embedding for each concept, and thus for each node. To prevent information leakage, embeddings used for training and testing are computed only from text available up to the corresponding cutoff year.
Link prediction
The previous method used by Krenn et al. to predict new links in their concept graph exclusively relied on abstract local graph properties, either through a purely graph-theoretical approach using hand-crafted features in conjunction with machine learning or through employing end-to-end machine learning methods [12, 14]. Here, we investigate whether integrating semantic knowledge about the concepts can improve link prediction. In particular, we use concept embeddings, i.e., high-dimensional vectors that capture semantic information, to make the semantic information integrable into the link prediction task.
Given the concept graph $G$, we treat link prediction as a binary classification task. The objective of the ML model is thus to predict whether a new edge is formed between an arbitrary pair of previously unconnected vertices $(u,v)$ in the time range $T=[T_{\mathrm{start}},T_{\mathrm{end}}]$. We chose $T_{\mathrm{start,train}}=2017$ and $T_{\mathrm{end,train}}=2019$ for training, i.e. our model has access to the entire data up to and including 2016 while its predictions are made for the years 2017, 2018, and 2019. We illustrate link prediction on a rudimentary concept graph in 8.
Figure 8: Example for a concept graph derived from two articles published 2012 (blue) in and 2015 (orange) with one overlapping concept. Possible new edges are marked as dotted gray lines together with their predicted probability of formation in the years 2017 to 2019.
The prediction task has an inherent strong label imbalance since the likelihood of a randomly selected vertex pair $(u,v)$ forming a link throughout three years is extremely low. While there are for example 18.7 billion possible new edges that could form between 2017 and 2019, only 1.3 million new edges ($0.007\%$) can actually be observed during this period. To address this imbalance, we oversampled positive labels by using a fixed percentage (30%) of positive examples per batch in the training process. This oversampling during training shifts the trade-off between precision and recall in imbalanced tasks toward higher recall at the cost of losing precision, thus favoring the generation of larger sets of suggestions that may contain inspiring concept combinations over smaller sets in which some valuable ideas might not be included.
A modified version of Krenn et al.’s densely connected neural network [14], which purely relies on graph properties at different points in time, was used as the Baseline model. Specifically, the degree of a node $u$ ($\sum_{i=1}^{n}A_{u,i}$) and the sum of all 2-length paths from $u$ ($\sum_{i=1}^{n}A_{u,i}^{2}$) were calculated for different years in the range $t=[T_{\text{start,train}}-5,T_{\text{start,train}}-1]$, where $A_{t}$ denotes the binary adjacency matrix of $G_{t}$. These features were then concatenated for a given pair of nodes $(u,v)$ to result in a 20-dimensional baseline feature vector. In the second Concept embeddings (MatSciBERT) model, the concatenated concept embeddings of $u$ and $v$ were used instead as the (1,536-dimensional) feature vector for the neural network classifier, to test their information content and relevance for the link prediction task. We repeat the embeddings generation process with BERT, yielding a modified Concept embeddings (BERT) model. To explore another way of utilizing semantic information, we fine-tune the MatSciBERT model directly to predict the likelihood of two concepts becoming connected in our concept graph, thus, yielding the Pure Text Baseline model.
The Baseline model was then combined with the Concept embeddings model in two hybrid models, for the first of which (Combination of features), a concatenation of the feature vectors of the two previous models was used as the input. The hyperparameters of all neural networks were optimized using a comprehensive grid search varying the number of neurons per layer, the percentage of positive samples in each batch, the learning rate, and the dropout probability (see Supplementary Table 9 for a list of optimized hyperparameters).
The second hybrid model (Mixture of Baseline and Embeddings) uses a weighted output of the optimized Baseline and Concept embeddings (MatSciBERT) models where an optimal weighting of 3:2 was determined using hyperparameter optimization. We note that there are many other possible hybrid models beyond concatenating the two input vectors and calculating the weighted average of the output probabilities. The two parts of the input could be passed through a first set of layers separately before the two outputs are concatenated and passed through a second set of layers. The optimization of the architecture of the NN was however outside the scope of this study, as our goal mainly consisted in showing that the inclusion of the concept embeddings improves link prediction.
To explicitly capture local neighborhood structures through message passing, we implemented a graph neuralnetwork (GNN) model. Given the large-scale and hub-heavy nature of the network, we employed neighbor sampling to enable efficient training, initializing the node representations with the topological vectors from the Baseline model. This architecture utilizes a 2-layer GraphSAGE encoder [46] with neighbor sampling to compute node embeddings based on the graph topology at $T_{\text{start,train}}$. An MLP decoder was employed to classify the concatenated node embeddings.
Finally, we constructed a Mixture of GNN and Embeddings model as a third hybrid model, analogous to the Mixture of Baseline and Embeddings approach described above. This ensemble calculates a weighted average (1:1) of the output probabilities from the GNN and the Concept embeddings (MatSciBERT) models.
Human domain experts
The human domain experts who participated in the interviews are affiliated with different institutes at different institutions and were recruited to cover a wide range of topics within materials science. 13 researchers were invited to participate, 10 of which agreed to participate in the study and to conduct interviews. All participants are professors or independent group leaders. No interviews were excluded from the study.